OpenStudy (anonymous):

Stefan flips a coin four times. Which of the following shows the number of possible outcomes?

OpenStudy (anonymous):

2 8 sixteen sixty-four

OpenStudy (anonymous):

please help i am hopeless in math

OpenStudy (anonymous):

will anyone help me

OpenStudy (anonymous):

anybody.....anybody....*foreveralone*

OpenStudy (anonymous):

2 choices for each flip, 4 flips, all together there are \[2\times 2\times 2\times =2^4=16\]possible outcomes

OpenStudy (anonymous):

thanks so mush mate helped me my friend alot. but mostly my friend

OpenStudy (anonymous):

yw